OpenStudy (sonyalee77):

PLEZ HELP ME!!!!!!!!!! . Find the slope of the line. (-1,1) and (3,1)

OpenStudy (anonymous):

Slope = (3-1)/(1-(-1)) = 2/2 = 1

OpenStudy (sonyalee77):

thanx

OpenStudy (anonymous):

slope=Y2-Y1/X2-X1 =1-1/3-(-1) =0